# Universal and non-universal finite-volume effects in the vicinity of chiral phase transition in (2+1)-flavor QCD

### v1

@@ 1 Introduction
Despite having commendable lattice Quantum Chromodynamics (QCD) evidences for the restoration of spontaneously broken chiral symmetry, $SU(2)_{L}\times SU(2)_{R}$, in the limit of vanishing light quark masses through a second order phase transition [^1] For a recent overview and further references, see [17]. along with somewhat known broken nature of axial $U(1)_{A}$ symmetry at the chiral phase transition temperature $T_{c}$ [2, 4], the role of the axial anomaly in the vicinity of the chiral phase transition, in particular the strength of its breaking and its impact on physical QCD thermodynamic observables continues to remains to be an open question. It is therefore quite crucial to provide further concrete support for the universal critical behavior of chiral phase transition, which in the limit of two vanishing light quark masses does resemble the universality class of the $3$-dimensional, $O(4)$ spin model [16], and which is also supposed to provide stringent constraints on the location of the elusive QCD critical point in the QCD phase diagram [3, 7]. Although several studies, comparing lattice QCD calculations with expected universal critical behavior close to the chiral limit, are already existent in the related state-of-the-art literature, a majority of these numerical studies base their point for observation of universal critical behavior on the very presumption that the transition belongs to the $3$-d, $O(4)$ universality class and on subsequently constructing the corresponding universal scaling ansatz to parameterize deviations away from criticality in terms of additional non-singular terms contributing to the relevant observables. Thus, a direct lattice QCD determination of the critical parameters influencing the QCD chiral phase transition obviously is desirable.
@@ 4 Summary and Outlook
We presented first results from an ongoing study of finite-size effects in calculations of an improved order parameter for chiral symmetry breaking in $(2+1)$-flavor QCD, in which we attempt to quantify the finite-size effects as function of lattice volume and the working quark masses for a given temperature close to the chiral phase transition. Our calculations on lattices with fixed lattice cut-off corresponding to a temporal extent of $N_{\tau}=8$ show that precision studies of the order parameter require aspect ratios for the lattice size in units of temperature, $N_{\sigma}/N_{\tau}\geq 6$ at physical values of the quark masses, which increases to about $8$ for the currently used smallest quark mass ratio $H=m_{\ell}/m_{s}=1/240$. This is currently a work in progress.

### v2

@@ 1 Introduction
Despite having commendable lattice Quantum Chromodynamics (QCD) evidences for the restoration of spontaneously broken chiral symmetry, $SU(2)_{L}\times SU(2)_{R}$, in the limit of vanishing light quark masses through a second order phase transition [^1] For a recent overview and further references, see [17]. along with the explicitly broken nature of the $U(1)_{A}$ symmetry at the QCD transition temperature [2, 4], the role of the axial anomaly in the vicinity of the chiral phase transition, in particular the strength of its breaking and its impact on physical QCD thermodynamic observables continues to remains to be an open question. It is therefore quite crucial to provide further concrete support for the universal critical behavior of chiral phase transition, which in the limit of two vanishing light quark masses does resemble the universality class of the $3$-dimensional, $O(4)$ spin model [16], and which is also supposed to provide stringent constraints on the location of the elusive QCD critical point in the QCD phase diagram [3, 7]. Although several studies, comparing lattice QCD calculations with expected universal critical behavior close to the chiral limit, are already existent in the related state-of-the-art literature, a majority of these numerical studies base their point for observation of universal critical behavior on the very presumption that the transition belongs to the $3$-d, $O(4)$ universality class and on subsequently constructing the corresponding universal scaling ansatz to parameterize deviations away from criticality in terms of additional non-singular terms contributing to the relevant observables. Thus, a direct lattice QCD determination of the critical parameters influencing the QCD chiral phase transition obviously is desirable.
@@ 4 Summary and Outlook
We presented first results from an ongoing study of finite-size effects in calculations of an improved order parameter for chiral symmetry breaking in $(2+1)$-flavor QCD, in which we attempt to quantify the finite-size effects as function of lattice volume and the light quark masses for a given temperature close to the chiral phase transition. Our calculations on lattices with fixed lattice cut-off corresponding to a temporal extent of $N_{\tau}=8$ show that precision studies of the order parameter require aspect ratios for the lattice size in units of temperature, $N_{\sigma}/N_{\tau}\geq 6$ at physical values of the quark masses, which increases to about $8$ for the currently used smallest quark mass ratio $H=m_{\ell}/m_{s}=1/240$. This is currently a work in progress.